If the pressure of a gas contained in a closed vessels is increased by 0.4% when heated by 1-degree Celsius its Initial temperature must be -

Grade:11

1 Answers

Arun
6 years ago
Let P2 = 1.004 P1 ..... A 0.4% increase in P 
Let T2 = T1 + 1 ........... See note below 

P1V1/T1 = P2V2/T2 ..... combined gas law 
P1/T1 = P2/T2 ............... Assume V remains constant. 
T1 = P1T2 / P2 .............. solve for T1 
T1 = P1(T1+1) / (1.004P1) 
T1 = (T1+1) / 1.004 
1.004T1 = T1 + 1 
1.004T1 - T1 = 1 
0.004T1 = 1 
T1 = 1 / 0.004 
T1 = 250K 

Note: A 1 degree increase on the Celsius scale is a 1 degree increase on the Kelvin scale. Temperatures used in the combined gas law must be in Kelvin.